# mmCooper: A Multi-agent Multi-stage Communication-efficient and Collaboration-robust Cooperative Perception Framework

### v1

@@ 4.3 Ablation Study
GSP MOA FBCFormer LF IF AP@0.7/0.5()($\uparrow$)( ↑ )
@@
76.83/88.65 44.11/59.92
66.80/78.62 46.21/57.02
Table 3: Ablation study results of different designs in mmCooper on the OPV2V and DAIR-V2X datasets. GSP: Gumbel Softmax-based Partitioner; MOA: Multi-scale Offset-aware Attention; FBCFormer: Feature-guided Bounding Boxes Calibration Transformer; LF: Late-stage Fusion; IF: Intermediate-stage Fusion.
We conducted comprehensive ablation studies on the OPV2V and DAIR-V2X datasets to demonstrate the importance of different designs, with the ablation results shown in Table 3.
Importance of Core Components. All our core components contribute to performance improvement. Removing the Gumbel Softmax-based Partition (GSP) from the Confidence-based Filter Generation Module leads to a performance drop. This indicates that low-quality information exists in the intermediate-stage features and the late-stage bounding boxes, potentially due to irrelevant background or noise. Introducing adaptive and dynamic allocation for intermediate-stage features and late-stage bounding boxes effectively filters out irrelevant information. The absence of the intermediate-stage multi-scale Offset-aware Attention (MOA) module and the late-stage Feature-guided Bounding Boxes Calibration Transformer (FBCFormer) module also leads to a loss of accuracy, demonstrating that incorporating spatial neighborhood information and cross-stage calibration modules effectively addresses potential misalignment and calibration noise.
Results of Single-stage Fusion. We entirely removed the Late Fusion (LF) and the Intermediate Fusion (IF) separately to degrade our model to single-stage fusion. The experimental results indicate that removing either module leads to a performance decline, underscoring the effectiveness of the proposed multi-stage cooperation approach.
@@ 4.4 Qualitative Evaluation
Visualization of Detection Results. To visually display the detection results of our proposed mmCooper and the other baseline models, Figure 6 shows the visualization results for two scenarios in the DAIR-V2X dataset. Compared to SOTAs, our mmCooper achieves high-precision 3D object detection in these challenging scenarios, almost correctly predicting all ground truth. Specifically, the advantages of our method stem from two aspects: 1) The proposed dynamic adaptive allocation strategy enables dynamic partitioning of intermediate-stage features and late-stage bounding boxes, effectively filtering out irrelevant and noisy information. 2) Our Multi-scale Offset-aware Attention Module and BBox Filtering & Calibration Module effectively address misalignment and calibration noise.
@@ 1 Overview
  • Sec. 3: Additional Experimental Results on OPV2V and DAIR-V2X Datasets
@@ 2 The System Pipeline of mmCooper
The proposed system pipeline of mmCooper is illustrated in Alg.1. Note that the following pipeline is executed in parallel across all agents. For simplicity, we describe the pipeline from the perspective of the ego agent. The ego agent is represented by $i$, while the collaborative agents are denoted by $j$. First, the agent generates BEV features $F_{bev}^{i}$ and $F_{bev}^{j}$ through the Observation Module. In the Broadcasting Module, agents generate initial coarse bounding boxes $B_{coarse}^{i}$ and $B_{coarse}^{j}$, where $b^{i}$ and $b^{j}$ represent the number of bounding boxes predicted by the ego agent and the collaborating agents, respectively. The collaborative agents then package and broadcast the filtered BEV features along with the coarse bounding boxes. Specifically, $f_{gaussian}(\cdot)$ denotes a Gaussian filter, and $\{F_{bev,m}^{j},B_{m}^{j}\}$ represents the filtered features and bounding boxes from the collaborative agents.
Subsequently, in the Intermediate-stage Fusion, the ego agent performs feature fusion using the Multi-scale Offset-aware Attention module. The outputs from the fused features at different scales, after undergoing upsampling, are concatenated to obtain the final fused features. Here, $f_{up2}(\cdot)$ and $f_{up3}(\cdot)$ denote the upsampling operations, while $\{F_{i}^{sc1},F_{i}^{sc2},F_{i}^{sc3}\}$ represent the feature fusion outputs at three different scales.
@@
Finally, the fused features are input into a detection head to predict the fused bounding box $\mathcal{B}_{fused}$. The final bounding box is composed of the initial prediction, the bounding box predicted from the fused features, and the bounding box obtained after filtering and calibration from collaborative agents. $f_{post}(\cdot)$ represents the combination of all bounding boxes.
@@
$X_{j}^{{}^{\prime}}=T_{j->i}\cdot X_{j}$
$F_{bev}^{j}=\psi_{encoder}(X_{j}^{{}^{\prime}})\in\mathbb{R}^{C\times H\times W}$
@@
$B_{coarse}^{j}=f_{dec}(F_{bev}^{j})\in\mathbb{R}^{b^{j}\times 7}$
$C_{f}^{j},C_{b}^{j}=\phi_{conf}\left(F_{bev}^{j}\right)$
@@
$\mathcal{M}_{f/b}^{j}=M_{f/b,g}^{j}\odot M_{f/b,top}^{j}$
$F_{bev,m}^{j}=\mathcal{M}_{f}^{j}\odot F_{bev}^{j}$; $B_{m}^{j}=\mathcal{M}_{b}^{j}\odot B_{coarse}^{j}$
broadcast $\{F_{bev,m}^{j},B_{m}^{j}\}$ to other agent
@@
# Observation Module.
$F_{bev}^{i}=\psi_{encoder}(X_{i})\in\mathbb{R}^{C\times H\times W}$
# Broadcasting Module.
$B_{coarse}^{i}=f_{dec}(F_{bev}^{i})\in\mathbb{R}^{b^{i}\times 7}$
@@
Receive $\{F_{bev,m}^{j},B_{m}^{j}\}$ sent by collaborative agents.
Encode the feature set $\{F_{bev}^{i},F_{bev,m}^{j}\}$ into three scales $S=\{sc1,sc2,sc3\}.$
@@
$\mathcal{F}_{i,c}^{sc}=CrossAttn(\sigma(F_{bev,c}^{i,sc}),F_{bev,c}^{j,s\times s,sc},F_{bev,c}^{j,s\times s,sc})$
@@
$\mathcal{F}_{i}=concat(\mathcal{F}_{i}^{sc1},f_{up2}(\mathcal{F}_{i}^{sc2}),f_{up3}(\mathcal{F}_{i}^{sc3}))$
@@
$F_{b}(q)=f_{enc,b}(B_{i,m})$
$DBA(q)=\sum_{\alpha=1}^{A}W_{\alpha}[\sum_{n=1}^{N}\sum_{m=1}^{M}\omega(W_{\beta}F_{b}(q))F_{j}(q+\bigtriangleup q_{m})]$
$F_{DBA}=DBA(q)+\mathcal{F}_{i}$
@@
$\mathcal{B}_{m}^{j}=\varphi(off,score,B_{m}^{j})$
@@
$\mathcal{B}_{fused}=f_{dec}(\mathcal{F}_{i})$
$\mathcal{B}_{final}^{i}=f_{post}(B_{coarse}^{i},\mathcal{B}_{m}^{j},\mathcal{B}_{fused})$
3 Additional Experimental Results on OPV2V and DAIR-V2X Datasets
In this section, we provide additional experiments to supplement the results on the datasets OPV2V[43] and DAIR-V2X[49] presented in the main paper.
@@ 3.1 Implementation Details
On the OPV2V and DAIR-V2X datasets, the dimensions of the voxels encoded by the encoder are $0.4\times 0.4\times 4$. The shape of the shared BEV features among agents is (64, 100, 252) for DAIR-V2X and (64, 100, 352) for OPV2V. The shared bounding boxes among agents are represented by their center coordinates, dimensions (length, width, height), and heading angle. The detection decoder consists of two distinct 1×1 convolutional layers.
@@ 3.2 Supplements on Localization Errors
The localization errors on the OPV2V and DAIR-V2X datasets are sampled from a Gaussian distribution with a mean of 0 $m$ and a standard deviation $\sigma\in\{0.0,0.1,0.2,0.3,0.4\}m$. The experimental results, as shown in Figure 7, demonstrate that our proposed mmCooper outperforms the existing state-of-the-art methods[23, 11, 18, 42, 50] across different levels of localization error, highlighting its robustness to such errors.
@@ 3.3 Supplements on Transmission Delays
We evaluated the performance of the models on the OPV2V and DAIR-V2X datasets under transmission delays for $\{0,100,200,300,400\}ms$. As shown in Figure 8, our proposed mmCooper consistently outperforms the existing state-of-the-art methods across all transmission delay conditions, demonstrating the superiority of our approach in scenarios with transmission delays.
Figure 7: Robustness to the localization error on the OPV2V and DAIR-V2X datasets.
Figure 8: Robustness to the transmission delay on the OPV2V and DAIR-V2X datasets.
Figure 9: Robustness to the heading error on the OPV2V and DAIR-V2X datasets.
dui
@@
Figure 9 demonstrates the performance of our proposed mmCooper method compared to other baseline methods under varying levels (i.e., $\{0.0,0.2,0.4,0.6,0.8\}^{\circ}$) of heading error noise on the OPV2V and DAIR-V2X datasets. As illustrated in Figure 9, the performance of all models decreases as the heading errors increase. However, mmCooper consistently outperforms the current state-of-the-art models across all error levels, highlighting the advantages of our designed Multi-Scale Offset-Aware Fusion Module and BBox Filtering & Calibration Module in enhancing system robustness.
@@ 4 Additional Qualitative Evaluation Results
We provide additional qualitative results on DAIR-V2X datasets. Figures 10 present visualizations of different road scenarios. Our proposed mmCooper method can detect almost all ground truths without any false positives. The results demonstrate that our proposed mmCooper achieves outstanding detection performance across various scenarios.

### v3

@@ 4.3 Ablation Study
Table 3: Ablation study results of different designs in mmCooper on the OPV2V and DAIR-V2X datasets. CFG: Confidence-based Filter Generation Module; MOF: Multi-scale Offset-aware Fusion; BFC: BBox Filtering & Calibration Module; LF: Late-stage Fusion; IF: Intermediate-stage Fusion.
CFG MOF BFC LF IF AP@0.7/0.5()($\uparrow$)
@@
76.83/88.65 46.65/62.84
66.80/78.62 45.78/56.88
We conducted comprehensive ablation studies on the OPV2V and DAIR-V2X datasets to demonstrate the importance of different components, as shown in Tab. 3.
Importance of Core Components. All the core components contribute to performance improvement. Removing the CFG module results in the transmission of information for both the intermediate and late stages across all locations without suppression or selection. While this variant increases the amount of transmitted information compared to mmCooper, it leads to a performance drop. This highlights the importance of filtering out low-confidence information and selectively transmitting data at different stages to optimize performance. The absence of the MOF module replaces the MOA module with a self-attention mechanism, while removing the BFC module eliminates the filtering and refinement of received bounding boxes. Both modifications result in a performance drop, highlighting the importance of incorporating spatial neighborhood information to mitigate misalignment and applying bounding box filtering and refinement to address redundancy, imprecision, or potential errors in received detections.
Superiority of Multi-stage Fusion. We entirely removed the Late Fusion (LF) or the Intermediate Fusion (IF) to degrade our model to single-stage fusion. The experimental results indicate that removing either module leads to a performance decline, underscoring the effectiveness of the proposed multi-stage cooperation framework.
Figure 7: Visualization of two-stage fusion on DAIR-V2X. We show the results from the Confidence-based Filtering Generation Module, including discarded information (white background), BBoxes for transmission (red dots), and features for transmission (yellow background). We also show the BFC module results, including uncalibrated BBoxes (blue boxes), calibrated BBoxes (red boxes), and ground truth BBoxes (green boxes).
@@ 4.4 Qualitative Evaluation
Visualization of Two-stage Fusion. As shown in the Fig. 7, we visualize the transmission decisions and bounding boxes refinement in the two-stage fusion on the DAIR-V2X dataset. The yellow background and red dots show the selective transmission of data through two stages. The white background shows the locations discarded by the CFG module, showing a substantial amount of low-confidence or irrelevant information in the scenario. The removal and correction of the uncalibrated blue boxes demonstrates the BFC module’s effectiveness in filtering and refining received bounding boxes by eliminating redundancies and improving alignment with the ground truth.
Visualization of Detection Results. Fig. 8 shows visualization results for a scenario in the DAIR-V2X dataset. Compared to baseline methods, mmCooper demonstrates high-precision 3D object detection, accurately predicting nearly all ground truth objects. In contrast, baseline methods either fail to detect certain vehicles or produce inaccurate bounding boxes. †Refer to the supplementary materials for additional visualized detection results.
@@ 5 Conclusion
Acknowledgements. This work is supported by the National Natural Science Foundation of China under Grant 62272357 and 62302326, Wuhan Science and Technology Project for Key Research and Development under Grant 2024050702030090 and Wuhan Science and Technology Joint Project for Building a Strong Transportation Country under Grant 2024-2-7.
[…]
@@ 1 Overview
  • Sec. 3: Details of Multi-stage Fusion Method
  • Sec. 4: Additional Experimental Results
@@
    • Sec. 4.5: More Ablation on V2VSet Dataset
    • Sec. 4.6: Performance on V2V4Real Dataset
    • Sec. 4.7: Computation Costs
    • Sec. 4.8: Ablation of Deformable BBox Attention (DBA)
    • Sec. 4.9: Impact of Varying the Number of Agents
@@
    • Sec. 5.1: Visualization of Detection Result
    • Sec. 5.2: Visualization of Multi-stage Fusion
@@ 2 The System Pipeline of mmCooper
The proposed system pipeline of mmCooper is illustrated in Algorithm 1. Note that the following pipeline is executed in parallel across all agents. For simplicity, we describe the pipeline from the perspective of the ego agent. The ego agent is represented by $i$, while the collaborative agents are denoted by $j$. First, the agent generates BEV features $F^{i}$ and $F^{j}$ through the Observation Encoding. In the Information Broadcasting, agents generate initial coarse bounding boxes $B^{i}$ and $B^{j}$, where $N_{b}^{i}$ and $N_{b}^{j}$ represent the number of bounding boxes predicted by the ego agent and the collaborating agents, respectively. The collaborative agents then package and broadcast the filtered BEV features along with the coarse bounding boxes. Specifically, $f_{gaussian}(\cdot)$ denotes a Gaussian filter, and $\{\hat{F}^{j},\hat{\mathcal{B}}^{j}\}$ represents the filtered features and bounding boxes from the collaborative agents.
Subsequently, in the Intermediate-stage Fusion, the ego agent performs feature fusion using the Multi-scale Offset-aware Attention module. The outputs from the fused features at different scales, after undergoing upsampling, are concatenated to obtain the final fused features. Here, $f_{up2}(\cdot)$ and $f_{up3}(\cdot)$ denote the upsampling operations, while $\{F_{sc1}^{i},F_{sc2}^{i},F_{sc3}^{i}\}$ represent the feature fusion outputs at three different scales.
@@
Finally, the fused features are input into a detection head to predict the fused bounding box $\mathcal{B}_{fused}$. The final bounding box is composed of the bounding box predicted from the fused features $\mathcal{B}_{fused}$ and the bounding box obtained after filtering and calibration $\hat{\mathcal{B}}_{m}^{j}$from collaborative agents. $f_{post}(\cdot)$ represents the combination of all bounding boxes.
@@
$F^{j}=\psi_{encoder}(x^{j})\in\mathbb{R}^{C\times H\times W}$
@@
$\mathcal{B}^{j}=f_{dec}(F^{j})\in\mathbb{R}^{N_{b}^{j}\times 7}$
$C_{f}^{j},C_{b}^{j}=\phi_{conf}\left(F^{j}\right)$
@@
$\mathcal{M}_{f/b}^{j}=M_{f/b,g}^{j}\odot M_{f/b,t}^{j}$
$\hat{F}^{j}=\mathcal{M}_{f}^{j}\odot F^{j}$; $\hat{\mathcal{B}}^{j}=\mathcal{M}_{b}^{j}\odot\mathcal{B}^{j}$
broadcast $\{\hat{F}^{j},\hat{B}^{j}\}$ to other agent
@@
# Observation Encoding.
$F^{i}=\psi_{encoder}(x^{i})\in\mathbb{R}^{C\times H\times W}$
# Information Broadcasting.
$\mathcal{B}^{i}=f_{dec}(F^{i})\in\mathbb{R}^{N_{b}^{i}\times 7}$
@@
Receive $\{\hat{F}^{j},\hat{\mathcal{B}}^{j}\}$ sent by collaborative agents.
Encode the feature set $\{F^{i},\hat{F}^{j}\}$ into three scales $S=\{sc1,sc2,sc3\}.$
@@
$\mathcal{F}_{sc,(r,c)}^{i}=CrossAttn(MLP(F_{sc,(r,c)}^{i}),\hat{F}_{sc,(r,c)}^{j,s\times s},\hat{F}_{sc,(r,c)}^{j,s\times s})$
@@
$\mathcal{F}^{i}=concat(\mathcal{F}_{sc1}^{i},f_{up2}(\mathcal{F}_{sc2}^{i}),f_{up3}(\mathcal{F}_{sc3}^{i}))$
@@
$F_{b}(q)=f_{enc,b}(\hat{\mathcal{B}}^{j})$
$DBA(q)=\sum_{\alpha=1}^{A}W_{\alpha}[\sum_{n=1}^{N}\sum_{m=1}^{M}\omega(W_{\beta}F_{b}(q))\mathcal{F}^{i}(q+\bigtriangleup q_{m})]+F_{b}(q)$
@@
$\hat{\mathcal{B}}_{m}^{j}=\varphi(off,score,\hat{\mathcal{B}}^{j})$
@@
$\mathcal{B}_{fused}=f_{dec}(\mathcal{F}^{i})$
$\mathcal{B}_{final}^{i}=f_{post}(\mathcal{B}^{i},\hat{\mathcal{B}}_{m}^{j},\mathcal{B}_{fused})$
3 Details of Multi-stage Fusion Method
Note that our proposed model requires only a single round of processing and communication, the same as other single-stage methods. As shown in Fig. 9, both intermediate-stage features and late-stage bboxes are handled within a single inference cycle and jointly transmitted in one communication round. While mmCooper includes two detection heads, they share parameters, and each head introduces only 2.71ms of processing time. Notably, mmCooper is more efficient than ERMVP, a single-stage intermediate fusion method, with a lower total runtime of 45.41ms compared to ERMVP’s 55.30ms. Combined with its superior performance, this confirms that mmCooper remains both practically feasible and system-efficient, without the overhead suggested in the comment.
Figure 9: Comparison of single-stage method and our method.
4 Additional Experimental Results
In this section, we provide additional experiments to supplement the results on the datasets OPV2V [50], DAIR-V2X [57], V2XSet [49] and V2V4Real [52] presented in the main paper.
@@ 4.1 Implementation Details
On the OPV2V, DAIR-V2X and V2XSet datasets, the dimensions of the voxels encoded by the encoder are $0.4\times 0.4\times 4$. The shape of the shared BEV features among agents is (64, 100, 252) for DAIR-V2X and (64, 100, 352) for OPV2V and V2XSet. The shared bounding boxes among agents are represented by their center coordinates, dimensions (length, width, height), and heading angle. The detection decoder consists of two distinct 1×1 convolutional layers.
@@ 4.2 Supplements on Localization Errors
The localization errors on the OPV2V, DAIR-V2X and V2XSet datasets are sampled from a Gaussian distribution with a mean of 0 $m$ and a standard deviation $\sigma\in\{0.0,0.1,0.2,0.3,0.4\}m$. The experimental results, as shown in Fig. 10, demonstrate that our proposed mmCooper outperforms the existing state-of-the-art methods [29, 13, 22, 49, 59] across different levels of localization error, highlighting its robustness to such errors.
@@ 4.3 Supplements on Transmission Delays
We evaluated the performance of the models on the OPV2V, DAIR-V2X and V2XSet datasets under transmission delays for $\{0,100,200,300,400\}ms$. As shown in Fig. 11, our proposed mmCooper consistently outperforms the existing state-of-the-art methods across all transmission delay conditions, demonstrating the superiority of our approach in scenarios with transmission delays.
@@
Fig. 12 demonstrates the performance of our proposed mmCooper method compared to other baseline methods under varying levels (i.e., $\{0.0,0.2,0.4,0.6,0.8\}^{\circ}$) of heading error noise on the OPV2V, DAIR-V2X and V2XSet datasets. As illustrated in Fig. 12, the performance of all models decreases as the heading errors increase. However, mmCooper consistently outperforms the current state-of-the-art models across all error levels, highlighting the advantages of our designed Multi-Scale Offset-Aware Fusion Module and BBox Filtering & Calibration Module in enhancing system robustness.
4.5 More Ablation on V2XSet Dataset
We supplement the main paper with ablation experiments conducted on the V2XSet dataset. As shown in Tab. 4, the results align with those presented in the main paper, demonstrating that the absence of any key component leads to performance degradation. Moreover, downgrading mmCooper to a single-stage approach also results in decreased performance.
Table 4: Ablation study results of different designs in mmCooper on the V2XSet datasets. CFG: Confidence-based Filter Generation Module; MOF: Multi-scale Offset-aware Fusion; BFC: BBox Filtering & Calibration Module; LF: Late-stage Fusion; IF: Intermediate-stage Fusion.
CFG MOF BFC LF IF AP@0.7/0.5()($\uparrow$)
V2XSet
65.86/84.40
Importance of Core Components
62.78/82.47
64.33/84.04
64.70/83.59
Results of Single-stage Fusion
64.19/83.73
54.36/72.14
Table 5: Collaborative perception performance on the V2V4Real dataset with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2° using Average Percision(AP)@0.7/0.5 as metrics.
Models V2V4Real
AP@0.7/0.5
No Fusion* [18] 24.72/43.52
Late Fusion* [18] 17.79/39.91
Intermediate Fusion* [18] 28.77/49.19
Where2comm [13] 29.83/49.43
V2X-ViT [49] 25.08/47.85
ERMVP [59] 26.81/44.57
Ours 31.47/50.21
4.6 Performance on V2V4Real Dataset
To further evaluate the performance of our model on more datasets, we also report the results on the V2V4Real dataset. Our experimental setup on the V2V4Real dataset is consistent with that of the DAIR-V2X dataset. As shown in Tab. 5, mmCooper still demonstrates outstanding performance on the V2V4Real dataset.
4.7 Computation Costs
Tab. 6 presents the overall computation time of our model as well as the computation time of each module. Although both our intermediate-stage and late-stage fusion modules utilize attention-based mechanisms, due to the sparsity of intermediate-stage features from collaborators and the sparsity of reference points used in deformable attention during late-stage fusion, our model achieves inference time comparable to other models.
4.8 Ablation of Deformable BBox Attention (DBA)
To further demonstrate the necessity of the Deformable BBox Attention (DBA) module, we replace it with the standard cross-attention mechanism. As shown in Tab. 7, due to the lack of focus on key reference points, replacing DBA with standard cross-attention leads to performance drops of 4.07%/0.07%, 0.58%/1.97%, and 1.44%/0.69% in AP@0.7/0.5 on OPV2V, DAIR-V2X, and V2XSet respectively.
Models OPV2V
Inference Time(ms)
Intermediate Fusion* 10.07
Where2comm [13] 17.05
V2X-ViT [49] 114.11
Select2col [28] 27.00
ERMVP [59] 55.30
Ours 45.41
Module OPV2V
Inference Time(ms)
Observation Encoding 12.26
Information Broadcasting 14.37
Intermediate-stage Fusion 13.54
Late-stage Fusion 4.06
mmCooper 45.41
Table 6: The left table presents the inference time of different models on the OPV2V dataset, while the right table shows the inference time of different components in mmCooper.
Table 7: Comparison of detection performance when using DBA or cross attention in the BBox Filtering & Calibration (BFC) module in AP@0.7/0.5 on OPV2V, DAIR-V2X, and V2XSet respectively.
Module OPV2V DAIR-V2X V2XSet
Cross Attn 74.04/88.86 47.69/63.15 64.42/83.71
DBA 78.11/88.93 48.27/65.12 65.86/84.40
4.9 Impact of Varying the Number of Agents
Fig. 13 illustrates the impact of varying the number of agents on detection performance in the OPV2V dataset. It can be observed that as the number of collaborating agents increases, the detection performance also improves.
@@ 5 Additional Qualitative Evaluation Results
5.1 Visualization of Detection Result
We provide additional qualitative results on DAIR-V2X datasets. Fig. 15 present visualizations of different road scenarios. Our proposed mmCooper method can detect almost all ground truths without any false positives. The results demonstrate that our proposed mmCooper achieves outstanding detection performance across various scenarios.
5.2 Visualization of Multi-stage Fusion
We provide more visualizations for multi-stage fusion in different scenarios. As shown in the Fig. 14, our mmCooper effectively performs dynamic allocation of intermediate-stage features and late-stage bounding boxes across different scenarios while also refining the received bounding boxes.
Figure 10: Robustness to the localization error on the OPV2V, DAIR-V2X and V2XSet datasets.
Figure 11: Robustness to the transmission delay on the OPV2V, DAIR-V2X and V2XSet datasets.
Figure 12: Robustness to the heading error on the OPV2V, DAIR-V2X and V2XSet datasets.
Figure 13: Impact of Varying the Number of Vehicle.
Figure 14: Visualization of two-stage fusion on DAIR-V2X. We show the results from the Confidence-based Filtering Generation Module, including discarded information (white background), BBoxes for transmission (red dots), and features for transmission (yellow background). We also show the BFC module results, including uncalibrated BBoxes (blue boxes), calibrated BBoxes (red boxes), and ground truth BBoxes (green boxes).